# Model averaging with mixed criteria for estimating high quantiles of extreme values: Application to heavy rainfall

### v1

@@ 6.2 Weighting scheme based on forward cross-validation
Contrary to the L-moments distance approach, many researchers have used cross-validation (CV) to determine weights for MA. Hansen and Racine (2012) proposed jacknife MA, where weights are selected based on leave-one-out CV. For time series MA, forward CV has been considered to preserve the temporal order in prediction (Hjorth 1982; Zhang & Zhang 2023). Although the data in this study are assumed to be stationary, we applied the forward CV approach to the weights selection, as our primary goal is to predict high quantiles, mainly out-of-sample quantities (i.e., extrapolative predictions).
@@
where the superscript `$te$’ denotes the test set, $n_{te}$ is the number of observations in the test set, and $s^{2}_{i,k}=\text{Var}(\hat{y}^{te}_{i,k})$ represents the variance estimate of the prediction, obtained in Section 3.4. Here, $\hat{y}^{te}_{i,k}$ is computed using the plotting position estimation, where the parameters are estimated from the training set. To compute $\hat{y}^{te}_{i,k}$, we consider the MLE and LME within each submodel $M_{k}$. A smaller value of $FCV_{k}$ indicates better extrapolative prediction performance, which is desirable when working with a limited number of extreme observations.To assign a higher weight to submodels with smaller $\text{FCV}_{k}$ values, we adopted the following likelihood function of the normal distribution:
@@
However, throughout our simulation study and real-data applications, we observed that the performance of the MA method with weights based on FCV is susceptible to the proportion of the training set (also referred to as the window size). We found that an appropriate proportion for reliable estimation depends on the return period and tail heaviness. In this study, we could not systematically determine an optimal proportion that simultaneously accounts for both the return period and tail heaviness. Nevertheless, we believe that the FCV is a promising criterion for weight assignment and should be further investigated in future research to enhance the accuracy of high-quantile predictions.
@@ 6.3 Extension to other models
Since it is known that the MA method generally performs well when the error variance is large (Liu et al. 2016, 2023), the proposed MA approach may apply to certain ‘sub-asymptotic’ extreme value distributions that suffer from high error variance in high quantile estimation. One example is the four-parameter kappa distribution (Hosking 1994; Shin & Park 2024; Strong et al. 2025), a generalization of the GEVD with one more shape parameter. Given that the MLE for the four-parameter kappa distribution exhibits high error variance (Papukdee et al. 2022), the MA method may provide a more reliable estimator than the MLE.
Extending the proposed method to non-stationary (NS) extreme value models would also be beneficial. In NS extreme value modeling, the MLE is highly sensitive to outliers or a few influential observations toward the end of the sample, often leading to substantial estimation variance for high quantiles. To address this issue, some researchers (Strupczewski & Kaczmarek 2001; Gado & Nguyen 2016) have proposed combining least squares and L-moment estimation methods. To fit an NS GEV model using the proposed MA approach, we first select $K$ submodels with fixed shape parameters. In NS GEV modeling, the shape parameter ($\xi_{t}$) is difficult to estimate precisely, so we treat $\xi$ as a constant. Under each fixed shape parameter $\xi_{k}$, we estimate the MLE of NS location and scale parameters. Next, the NS GEV submodel is transformed into a stationary extreme value model (Coles 2001). Under this transformed stationary model, the generalized L-moments distance is computed to assign weight to each submodel $M_{k}$. Further research is needed to explore this topic in greater detail.
@@ 7 Conclusion
We proposed a new MA method for estimating high quantiles of the GEVD. The proposed approach consists of selecting submodels based on the confidence interval of the shape parameter, estimating submodel parameters, and assigning weights to each submodel. Different criteria were employed for parameter estimation and weight assignment. Submodels were constructed using the MLE and LME, while weights were calculated using generalized L-moment distances and likelihood-based methods. This mixed approach, which leverages the strengths of both MLE and LME, is a key factor in improving accuracy when estimating high quantiles. Additionally, we quantified the uncertainty of the MA estimator for return levels under a random-weight setting. Furthermore, for further analysis, a surrogate model for the MA method was developed and applied to a real dataset—annual maximum daily precipitation in Hae-nam, Korea.
Based on our simulation study, the proposed MA method with the ‘gLd’ weighting scheme performed well for $-0.25<\xi\leq 0$, while the MA method with the ‘like’ weighting scheme was effective for $\xi\leq-0.25$. In general, without considering specific values of $\xi$ (i.e., tail heaviness), we recommend using the LME, the MA method with the ‘like’ weighting scheme, and the restricted MLE (or mixed estimator) considered by Morrison & Smith (2002) and Ailliot et al. (2011).
@@
We thank Prof. Sanghoo Yoon for his help for this study. This research was supported by Basic Science Research Program through the National Research Foundation of Korea(NRF) funded by the Ministry of Education(RS-2023-00248434), and by Global - Learning & Academic research institution for Master’s,PhD students, and Postdocs(LAMP) Program of National Research Foundation of Korea(NRF) grant funded by Ministry of Education(No. RS-2024-00442775)
@@
Yire Shin, $0$-0003-1297-5430;    Yonggwan Shin, $0$-0001-6966-6511
Jeong-Soo Park, $0$-0002-8460-4869
@@
Park conceived the study, Yonggwan Shin and Yire Shin conducted the analysis. All authors wrote the first draft, reviewed, edited, and approved the final manuscript.

### v2

@@ 7.2 Weighting scheme based on forward cross-validation
Contrary to the L-moments distance approach, many researchers have used cross-validation (CV) to determine weights for MA. Hansen and Racine (2012) proposed jackknife MA, where weights are selected based on leave-one-out CV. For time series MA, forward CV has been considered to preserve the temporal order in prediction (Hjorth 1982; Zhang & Zhang 2023). Although the data in this study are assumed to be stationary, we applied the forward CV approach to weight selection, as our primary goal is to predict high quantiles, mainly out-of-sample quantities (i.e., extrapolative predictions).
@@
where the superscript te`te’ denotes the test set, $n_{te}$ is the number of observations in the test set, and $s^{2}_{i,k}=\text{Var}(\hat{y}^{te}_{i,k})$ represents the variance estimate of the prediction, obtained in Section 3.4. Here, $\hat{y}^{te}_{i,k}$ is computed using the plotting position estimation, where the parameters are estimated from the training set. To compute $\hat{y}^{te}_{i,k}$, we consider the MLE and LME within each submodel $M_{k}$. A smaller value of $FCV_{k}$ indicates better extrapolative prediction performance, which is desirable when working with a limited number of extreme observations.To assign a higher weight to submodels with smaller $\text{FCV}_{k}$ values, we adopted the following likelihood function of the normal distribution:
@@
However, throughout our simulation study and real-data applications, we observed that the performance of the MA method with weights based on FCV is sensitive to the proportion of the training set (also referred to as the window size). We found that an appropriate proportion for reliable estimation depends on the return period and the tail heaviness. In this study, we could not systematically determine an optimal proportion that simultaneously accounts for both the return period and tail heaviness. Nevertheless, we believe that the FCV-based weighting is a promising criterion for weight assignment and should be further investigated in future research to enhance the accuracy of high-quantile predictions.
@@ 7.3 Extension to other models
Since it is known that the MA method generally performs well when the error variance is large (Liu et al. 2016, 2023), the proposed MA approach may be applicable to certain ‘sub-asymptotic’ extreme value distributions that suffer from high error variance in high quantile estimation. One example is the four-parameter kappa distribution (Hosking 1994; Shin & Park 2023; Strong et al. 2025), a generalization of the GEVD with an additional shape parameter. Given that the MLE for the four-parameter kappa distribution exhibits high estimation variance (Papukdee et al. 2022), the MA method may provide a more reliable estimator than the MLE.
Extending the proposed method to nonstationary (NS) extreme value models would also be beneficial. In NS extreme value modeling, the MLE is highly sensitive to outliers or a few influential observations toward the end of the sample, often leading to substantial estimation variance for high quantiles. To address this issue, some researchers (Strupczewski & Kaczmarek 2001; Gado & Nguyen 2016; Shin et al. 2025a) have proposed combining least squares and L-moment estimation methods. To fit an NS GEV model using the proposed MA approach, we first select $K$ submodels with fixed shape parameters. In NS GEV modeling, the shape parameter ($\xi_{t}$) is difficult to estimate precisely, so we treat $\xi_{t}$ as a constant. Under each fixed shape parameter $\xi_{k}$, we estimate the MLE of NS location and scale parameters. Next, the NS GEV submodel is transformed into a stationary extreme value model (Coles 2001). Under this transformed stationary model, the generalized L-moments distance is computed to assign weights to each submodel $M_{k}$. Further research is needed to explore this topic in greater detail.
@@ 8 Conclusion
We proposed a new MA method for estimating high quantiles of the GEVD. The proposed approach consists of selecting submodels based on the confidence interval of the shape parameter, estimating submodel parameters, and assigning weights to each submodel. Different criteria were employed for parameter estimation and weight assignment. Submodels were constructed using the MLE and LME, while weights were calculated using generalized L-moment distances and likelihood-based methods. This mixed approach, which leverages the strengths of both MLE and LME, is a key factor in improving accuracy in high quantile estimation. Additionally, we quantified the uncertainty of the MA estimator for return levels under a random-weight setting. Furthermore, a surrogate model for the MA method was developed for further analysis and applied to a real dataset–the annual maximum daily precipitation in Hae-nam, Korea.
Based on our simulation study, the proposed MA method with the ‘gLd’ weighting scheme performed well for $0.25<\xi\leq 0$, while the MA method with the ‘like’ weighting scheme was effective for $\xi\leq-0.25$. In general, without considering specific values of $\xi$ (i.e., tail heaviness), we recommend using the LME, the MA method with the ‘like1’ weighting scheme, and the restricted MLE (or mixed estimator) considered by Morrison & Smith (2002) and Ailliot et al. (2011). The Bayesian model averaging approach with a data-adaptive penalty function turned out to be a good alternative to existing estimation methods, especially for correcting the underestimation bias of the MA methods. We recommend using the BMA.like method as well.
@@
The authors are grateful to the reviewers and the associate editor of SERRA (Stoch Environ Res Risk Assess) for their valuable comments and constructive suggestions as these led to a greatly improved paper. We thank Prof. Sanghoo Yoon for his help for this study. This work was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2025-25436608).
@@
Yire Shin, 0000-0003-1297-5430;    Yonggwan Shin, 0000-0001-6966-6511
Jeong-Soo Park, 0000-0002-8460-4869
@@
Yire and Park conceived the study, Yonggwan and Yire conducted the analysis. All authors wrote the first draft, reviewed, edited, and approved the final manuscript.